3. The magnetic flux is a scalar that is proportional to the number of the magnetic field lines passing through a closed surface and is expressed by Φ) = BA cos θ, where B is the magnitude of the magnetic field, A the area of the surface and θ is the angle between the direction of the magnetic field and the line vertical to the surface. Calculate the magnetic flux through the area of the surface if the length of the rectangle is 8 m and the width is 5 m
1
Expert's answer
2020-04-21T10:57:50-0400

Area of the rectangle = 8*5 =40 m2

Flux = "BA\cos\theta=40B\cos\theta"

substituting the values of strength of magnetic field and cos"\theta" we will get the value of magnetic flux